1.0 Introduction

1.1 Spectrograph Characteristics

FAST has a conventional plane reflection grating geometry with an offaxis parabolic collimator and a Schmidt camera. The focal length of the collimator is 963 mm and the effective focal length of the camera is 328 mm, giving a reduction of 2.93, corresponding to a scale of 0.57 arcsec/15μ pixel along the slit. the unvignetted slit length is 3 arcminutes, or 13.72 mm at the slit. The angle between the collimated beams arriving at and leaving the grating is 35^{∘}. The layout of the spectrograph is shown in Figure 1.

1.2 Gratings

FAST is designed to accept gratings with rulings between 300 and 1200 lines per mm in first order. Table 1 gives the resolutions and spectral coverages for various gratings with a 40 mm long CCD, such as the Loral 512×2688. The spectrograph will vignette at the spectral extremes if a grating with an anamorphic factor exceeding 1.3 is used. This corresponds to a 1200 line grating centered at H_{α}, so that the use of a 1200 line grating in the extreme red will entail a modest loss in efficiency.

Ruling 2 Pixel Res Spectral Coverage 2 Pixel Slit Width
300 gpm  2.95 Å  3965 Å  1.21 arcsec
600 gpm  1.49 Å  2005 Å  1.25 arcsec
1200 gpm  0.75 Å  1005 Å  1.52 arcsec

We currently have three gratings available: a 300 and 600 line grating blazed at 4750 Å, and a 1200 line grating blazed at 5700 Å. The 300 and 600 line gratings have 50% efficiency limits of 3500 and 8000 Å, and the 1200 line grating has 5010000 Å. The gratings can be changed manually through the access hatch of the spectrograph.

1.3 Efficiency

FAST has been designed for high throughput between 3700 and 9000Å, making use of high efficiency antireflection coatings and UV enhanced, over coated silver coatings for the mirrors. With a back-side illuminated, thinned CCD with antireflection coating, the peak efficiency (including detector) should exceed 40%. With a thick, front side illuminated CCD, the peak efficiency should reach 20%. Figure 2 gives a plot of the estimated efficiency versus wavelength.

1.4 CCD Characteristics

The initial CCD is a Loral 512times2688 chip with 15μ pixels. The CCD is coated with a “KPNO Thick Green” phosphor coating to allow some blue efficiency. The CCD runs at about -112 ^{∘}C, giving a dark current of a few electrons/hr/pixel. The noise of the CCD is ∼8.5 electrons RMS. The CCD gain is currently set to 1.2 electrons per ADU. Be aware the the phosphor coating will have an afterglow after exposure to bright light, with a 1/e time constant of 15 minutes. It will take up to 2 hours for all effects of an overexposure to wear off.

We intend to replace the thick CCD with a thinned chip as soon as possible. This will double the quantum efficiency of the spectrograph and eliminate the phosphor afterglow.

1.5 Slits

There are five fixed width slits available: 1.1, 1.5, 2, 3, and 5 arcseconds wide. These can be changed manually through the access hatch of the spectrograph.

1.6 Blocking Filters

We have two long pass filters available with cutoffs of 3700 and 5000 Å. These can be mounted in a slide between the slit and the shutter.

1.7 Automated Features

The collimator focus and the comparison stage can be remotely operated via computer, as can the three calibration lamps (He-Ar, Ne, and Incand).

2.0 Calibration Lamps

The wavelength comparison lines are produced by a He-Ar lamp and a Ne lamp, which are ordinarily energized simultaneously. The relative intensities of the two lamps has been adjusted so that the required exposure times are identical. Figure 3 is a spectral line atlas over most of the useful spectral range with the 300 gpm grating and Figure 4 is an altlas in the UV and blue with the 1200 gpm grating to resolve some blended lines. Note the presence of some Fe lines in the UV; these are produced by the Fe cathode in the He/Ar lamp.

The incandescent lamp unfortunately is weak below 4500 Å, even after the introduction of a color balancing filter. Bluer filters are available, but the filters we tried introduced strong variations in intensity with periods of several hundred Å. We felt that most users would prefer to make multiple flat exposures rather than deal with strong short period intensity variations.

All three lamps are fed into an integrating sphere to scramble the lamp outputs and then pass through an achromatic lens to simulate the exit pupil of the telescope. This should produce an accurate calibration. Please be aware, however, that no calibration lamp is perfect, and that wavelength shifts between the sky and the lamps will be present at some level. A good way to check for offsets is to look at the calibrated wavelengths of known strong lines in the spectrum of the night sky.

All light from the integrating sphere passes through a 1 mm thick BG-14 filter. This knocks down the intensity of the strong red comparison lines and provides some color balancing for the incandescent lamp.

3.0 Focusing and Wavelength Stability

The spectrograph case is constructed from low-expansion graphite-epoxy composite panels with a coefficient of thermal expansion less than 25% that of aluminum. As a result, the spectrograph should stay in focus over the temperature changes that will be encountered on the typical run. The graphite epoxy composite will expand and contract with humidity changes, but this is a very slow process that should not be noticeable in a single run.

FAST was designed to achieve good short term stability, that is good stability over the course of the longest anticipated exposure, about one hour. There are two major contributors to this stability: mechanical flexure and thermal shifts. We will measure the mechanical flexure of the spectrograph during its commissioning. We have measured the thermal stability of the spectrograph in the laboratory and find a 0.2 pixel/^{∘}C shift, close to what we calculated should be present. Given that the night-time temperature changes are typically 0.25 ^{∘}/hour during the night, this should not be significant even for a long exposure. The thermal time constant of the spectrograph is about 3.5 hours, so if the dome is allowed to heat up significantly during the day, it would be wise to open the dome as soon as possible if the maximum stability is required.

Each step of the stepper motor that controls the collimator focus corresponds to a movement of 10 μ. Given that the telescope is f/10, that the reduction is 2.9 and the 15 μ pixel size, it will probably not be necessary to focus in moves smaller than 10–20 steps.

4.0 Changing Slits and Gratings

Please do not change slits or gratings without first receiving training from qualified personnel. We have secured the services of the FBI crime lab to identify all fingerprints in inappropriate places. Ex agents from an undisclosed foreign service (dismissed for excessive brutality) will take appropriate action to assure that damage to the equipment is not possible in the future. Please don’t misinterpret this friendly warning as any sort of threat.

The gratings and slit are installed through the single removal panel on the spectrograph. This panel is secured with two handles that rotate by 90removed and placed in a safe location before proceeding further. The slits held in place by spring loaded plungers, and the slit is simply pulled out by grasping its handle. The new slit should be inserted reflective side up and pushed home. The filter slide, mounted below the slit is removed and installed in the same fashion.

Changing gratings requires care. If you are not mechanically adept, please have Bas or Ted change the gratings for you during the day. Do not attempt to change gratings without a demonstration in any event! Grating replacements are between $2000 and $5000, and the lead time is long. Before removing the existing grating, please remove heavy jackets as the access port will not accommodate down jackets and heavy sweaters. The grating is secured with two bolts, one in back and one just inside the access panel. These two bolts must be loosened so that the grating rests on the slides that guide it into position. The arm that carries the forward bolt is then pressed up and rotated 90^{∘} to the right until the spring plunger seats to hold it out of the way. The grating can then be carefully removed, taking care not to place your fingers on or near the grating surface. As soon as the grating is in a safe location, the protective cover should be installed and the grating stowed. The cover for the new grating should be removed just before installation, and the removal procedure reversed. Tighten the two grating bolts gradually, alternating so that the gratin seats properly in its kinematic mount. Do not overtighten the bolts: modest finger tightening is sufficient!

5.0 Setting Grating Central Wavelengths

The grating tilt (and hence wavelength coverage) is adjusted with a manual, digital-readout micrometer. Each digit corresponds to a 0.001 inch adjustment. Two different styles of micrometers are used: the style used with the 300 and 600 gpm gratings have a locking ring, while the other style used for the 1200 gpm grating does not. We expect that the 1200 gpm grating micrometer has sufficient internal friction to prevent undesired rotation. Please alert us if you have any doubts on this score. The locking ring is loosened with a counterclockwise rotation. If the micrometer does not turn freely, do not force it. Please loosen the ring. The equations to calculate the correct micrometer setting are given below.

300 gpm grating:

Setting = 0.0771λ + 171

600 gpm grating:

Setting = 0.1488λ − 235

1200 gpm grating with no spacer (5400–7000Å:

Setting = 0.3053λ − 1663

1200 gpm grating with 0.45 inch spacer (4000–5500Å:

Setting = 0.2823λ − 1075

Please ask Bas to install the 0.45 inch spacer if you wish to use the 1200 gpm grating in the blue.

6.0 Scattered Light

Scattered light in the spectrograph results from undesired reflections from the refractive optics and the CCD, imperfections in the reflective surfaces and scattering of light outside first order from the spectrograph case. We have minimized these sources of scattered light as far as we know how, but some remains. Scattered light amounting to 3% of the dispersed light is scattered fairly uniformly across the CCD surface. Only about 1% of this scattered light (or 0.03% of the dispersed light) will fall on top of the desired spectrum, but strong sky lines may contribute as well.